Find (i) The total number of neutrons and (ii) The total mass of neutron in 7 mg of 14 C. (Assume that mass of neutron = mass of hydrogen atom)

Find
(i) The total number of neutrons and
(ii) The total mass of neutron in 7 mg of 14C.
(Assume that mass of neutron = mass of hydrogen atom)

Grade:10

1 Answers

Jitender Pal
9 years ago
(i) No. of C atom in 14g of 14C = 6.02 * 1023
∴ No. of C atom in 7 mg (7/1000g) of 14C
= 6.02 *1023 * 7/14 *1000 = 3.01 * 1020
No. of neutrons in 1 carbon atom = 7
∴ Total no. of neutrons in 7 mg of 14C = 3.01 * 1020 *7
= 21.07 *1020
Wt of neutron = wt of 1 hydrogen atom
= 1/6.02 *1023 g
∴ Wt of 3.01 * 1020 * 7 neutrons
= 3.0 *1020 *7/6.02 * 1023 = 3.5 * 10-3g